# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v1

@@ II.1 Recurrent neural network model
We study a recurrent neural network of $N$ neurons. Each neuron is characterized by its pre-activation $x_{i}(t)$ and activation $\phi_{i}(t)$, where $i=1,\ldots,N$. The activations are related to the pre-activations through a nonlinear function $\phi(x)$, with $\phi_{i}(t)=\phi(x_{i}(t))$. While our results are agnostic to the form of the nonlinearity, when comparing theory and simulations we use $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$, which is similar to the more conventional $\phi(x)=\tanh(x)$ but allows for analytical evaluation of Gaussian integrals. The network dynamics are governed by
$T[x_{i}](t)=\sum_{j}J_{ij}\phi_{j}(t),$ (1)
where $J_{ij}$ is the coupling from neuron $j$ to neuron $i$, and $T[\cdot]$ is a causal functional describing single-neuron dynamics. A canonical choice is
$T[x](t)=(1+\partial_{t})x(t),$ (2)
resulting in low-pass filtering of the input. We measure time in units of the time constant that would appear in front of $\partial_{t}$.
II.2 Correlation functions
II.2.1 Two-point functions
To characterize the network dynamics, we introduce several types of two-point correlation functions. Let ${a\in\{x,\phi\}}$ denote either the pre-activations or activations. We first define the neuron- and time-averaged covariance:
$C^{a}(\tau)=\frac{1}{N}\sum_{i}\left\langle a_{i}(t)a_{i}(t+\tau)\right\rangle_{t}.$ (3)
We also define the time-averaged, but not neuron-averaged, covariance; and the neuron-averaged, but not time-averaged, covariance:
@@ II.2.1 Two-point functions
$C^{a}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i}a_{i}(t_{1})a_{i}(t_{2}).$ (5)
For a given $\tau$, the spectra of $C^{a}_{ij}(\tau)$, with indices $i$ and $j$, and of $C^{a}(t_{1},t_{2}+\tau)$, with indices $t_{1}$ and $t_{2}$, are the same (up to a scaling), so spectral properties of one parameter can be computed using the other, a form of duality. For example, the properly scaled trace operations applied to these parameters yield the same result, namely, $C^{a}(\tau)$.
@@
We also define the response function:
$S^{a}_{ij}(t_{1},t_{2})=\left.\frac{\delta a_{i}(t_{1})}{\delta I_{j}(t_{2})}\right|_{I_{j}=0},$ (6)
where $I_{j}(t)$ is an infinitesimal source term added to the right-hand side of the network dynamics (Eq. 1). In analogy with the covariance functions, we define $S^{a}_{ij}(\tau)$, $S^{a}(t_{1},t_{2})$, and $S^{a}(\tau)$ to be the response function averaged over time, neurons, or both.
II.2.2 Four-point functions
Figure 2: Relationship between principal component timescales and two- and four-point correlation functions. (a) Example activity traces of principal components $p^{\phi}_{k}(t)$. (b) Empirical autocovariance functions for each principal component, weighted by their eigenvalues. (c) Same as (b) but weighted by the squared eigenvalue. (d) Averages over curves in (b) and (c), with comparisons to $C^{\phi}_{\star}(\tau)$ and $\Psi^{\phi}_{\star}(\tau,0)$. Dots: simulation results. Lines: theoretical predictions. All panels describe the same network of $N=4000$ neurons with i.i.d. couplings and $g=6$.
To characterize the dimension and timescales of activity, we define
$\Psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i,j}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}),$ (7)
where $\bm{\tau}=(\tau_{1},\tau_{2})$. This sum includes the on-diagonal contributions, unlike the quantity computed in Clark et al. [25], which we denote by the lower-case $\psi^{a}(\bm{\tau})$:
$\psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i\neq j}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}).$ (8)
For the type of connectivity we will study, the network exhibits high-dimensional activity such that pairwise covariances are $\mathcal{O}(1/\sqrt{N})$. Thus, in the definition of $\Psi^{a}(\bm{\tau})$ (Eq. 7), the on-diagonal contributions (given by $N^{-1}\sum_{i}C^{a}_{ii}(\tau_{1})C^{a}_{ii}(\tau_{2})$) and off-diagonal contributions (given by $\psi^{a}(\bm{\tau})$) are both $\mathcal{O}(1)$.
@@ II.2.2 Four-point functions
Leveraging neuron-time duality, $\Psi^{a}(\bm{\tau})$ can also be expressed in terms of the time-by-time covariance $C^{a}(t_{1},t_{2})$:
@@
The neuron-by-neuron (Eq. 7) and time-by-time (Eq. 9) definitions of $\Psi^{a}(\bm{\tau})$ form the basis of the cavity and path-integral calculations of this function, respectively.
II.2.3 Relating $\Psi^{a}$ to the dimension and timescales of activity
The dimension of activity can be quantified by the participation ratio of its covariance spectrum [34, 35, 23, 36, 32, 37, 9, 38, 27]. To define the participation ratio, we diagonalize the equal-time covariance matrix $C_{ij}^{a}(0)$ to obtain eigenvectors $\bm{v}^{a}_{k}$ and eigenvalues $\lambda^{a}_{k}$, as in principal components analysis. The participation ratio is
$\text{PR}^{a}=\frac{1}{N}\frac{(\sum_{k}\lambda^{a}_{k})^{2}}{\sum_{k}(\lambda^{a}_{k})^{2}},$ (10)
[…]
An advantageous property of this dimension measure is that it can be expressed in terms of the previously defined two- and four-point correlation functions as
$\text{PR}^{a}=\frac{C^{a}(0)^{2}}{\Psi^{a}(0,0)}.$ (11)
Recall that for the types of connectivity we study, cross-covariances are $\mathcal{O}(1/\sqrt{N})$, so the on- and off-diagonal contributions to $\Psi^{a}(\bm{\tau})$ are both order-one (Eq. 7). This leads to $\text{PR}^{a}$ taking a value between 0 and 1 as $N\rightarrow\infty$. Thus, the unnormalized dimension of activity $N\,\text{PR}^{a}$ is extensive, i.e., activity is high-dimensional. At the same time, activity does not uniformly occupy all available dimensions, which would require cross-covariances to be smaller than $\mathcal{O}(1/\sqrt{N})$.
Timescales of collective activity can be studied as follows. We define the principal components (PCs) of network activity as
$p^{a}_{k}(t)=\frac{(\bm{v}^{a}_{k})^{T}\bm{a}(t)}{\sqrt{\lambda^{a}_{k}}}.$ (12)
These PCs are the basis of much of modern analysis of high-dimensional neural data. The PCs all have unit variance but potentially very different characteristic timescales (Fig. 2a). To extract the timescales of just the leading PCs, we compute the PC-averaged covariance, weighted by the squares of the eigenvalues of the corresponding PCs. This gives:

### v2

@@ III.1 Recurrent neural network model
We study a recurrent neural network of $N$ neurons. Each neuron $i\in\{1,\ldots,N\}$ is characterized by its pre-activation $x_{i}(t)$ and activation $\phi_{i}(t)=\phi(x_{i}(t))$, where $\phi(\cdot)$ is a scalar nonlinearity; we use $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$  [^4] This behaves similarly to the more conventional $\phi(x)=\tanh(x)$ but allows for analytical evaluation of Gaussian integrals.. The network dynamics are governed by
$T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t),\>\>T[x](t)=(1+\partial_{t})x(t),$ (15)
where $J_{ij}$ denotes the synaptic coupling from neuron $j$ to neuron $i$, and $T[\cdot]$ is a causal functional that specifies the single-neuron dynamics. We expect our results to be agnostic to the specific choice of $T[\cdot]$. We next introduce correlation functions that capture single-neuron and collective activity properties.
III.2 Two-point functions and duality of neuronal and temporal covariances
Figure 3: Duality of neuron-by-neuron and time-by-time covariances and its relation to the dimension of activity. Both plots are based on the same simulation of a network of $N=2500$ neurons with $g=2.25$ and i.i.d. couplings. The dimension can be computed either by computing the statistics of the off-diagonals of the neuron-by-neuron covariance $C^{\phi}_{ij}(0)$ (left), or by computing the fluctuations away from the translation-invariant mean-field form of the time-by-time covariance $C^{\phi}(t_{1},t_{2})$ (right).
Let $a\in\{x,\phi\}$ denote either pre-activations or activations. We define two complementary covariance functions, averaging over either time or neurons:
@@ III.2 Two-point functions and duality of neuronal and temporal covariances
$C^{a}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}a_{i}(t_{1})a_{i}(t_{2}).$ (17)
These functions exhibit a form of duality: for any lag $\tau$, the matrix $\bm{C}^{a}(\tau)$ with elements $C^{a}_{ij}(\tau)$ indexed by $(i,j)$ and the matrix with elements $C^{a}(t_{1},t_{2}+\tau)$ indexed by $(t_{1},t_{2})$ have the same eigenvalue spectrum up to scaling (to see this, note that both matrices can be computed from a time-by-neuron activity matrix $\bm{A}$ as $\bm{A}^{T}\bm{A}$ and $\bm{A}\bm{A}^{T}$, respectively, which have identical spectra)  [^5] This duality is reminiscent of ergodicity theorems, which relate temporal and spatial averages, but differs in a variety of ways, including that it holds even for finite time intervals, whereas ergodicity concerns infinite-time limits..
@@
We also define the neuron-averaged response function, describing propagation of infinitesimal perturbations:
$S^{a}(t_{1},t_{2})=\frac{1}{N}\sum_{i=1}^{N}\left.\frac{\delta a_{i}(t_{1})}{\delta I_{j}(t_{2})}\right|_{I_{j}=0},$ (18)
where $I_{j}(t)$ is a source term added to the right-hand side of the network dynamics: $T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t)+I_{i}(t)$. Finally, we defined “stationarized” counterparts:
$C^{a}(\tau)=\left\langle C^{a}(t,t\pm\tau)\right\rangle_{t},\>\>\>\>S^{a}(\tau)=\left\langle S^{a}(t,t-\tau)\right\rangle_{t}.$ (19)
III.3 Four-point functions, dimension of activity, and principal-component timescales
In contrast to the above two-point functions, which characterize single-neuron activity, four-point functions describe features of collective activity. The same four-point functions were used in [30]. Here, we motivate these functions in more detail and express them in two ways via the duality between neuronal and temporal covariances.
To motivate their definition, consider the dimension of activity, which can be quantified as the participation ratio of the spectrum of the equal-time covariance matrix $\bm{C}^{a}(0)$ [44, 45, 29, 46, 47, 48, 9, 49, 50]. Given the eigenvectors $\bm{v}^{a}_{k}$ and eigenvalues $\lambda^{a}_{k}$ for $k=1,\ldots,N$ (as one would compute as part of principal components analysis), the participation ratio is
$\text{PR}^{a}=\frac{1}{N}\frac{(\sum_{k=1}^{N}\lambda^{a}_{k})^{2}}{\sum_{k=1}^{N}(\lambda^{a}_{k})^{2}}.$ (20)
To see why $\text{PR}^{a}$ provides a meaningful measure of dimension, consider the case where $D$ eigenvalues equal a positive constant and the remaining eigenvalues are zero. Then, $\text{PR}^{a}=D/N$. More generally, when the spectrum exhibits a smooth decay rather than a hard cutoff, the participation ratio identifies the characteristic decay scale (divided by $N$) [46]. We compare the participation ratio to alternative measures of effective dimensionality in Appendix D. Without the normalization factor $1/N$, this quantity would vary between $1$ and $N$; with normalization, it varies between $1/N$ and $1$.
In the limit $N\to\infty$, this quantity could have three qualitatively different behaviors:
  • Subextensive dimensionality: $\text{PR}^{a}=0$. The number of dimensions filled by activity grows sublinearly with $N$.
  • Nontrivial extensive dimensionality: $0<\text{PR}^{a}<1$. The number of dimensions filled by activity grows linearly with $N$ and not all dimensions are filled equally.
  • Trivial extensive dimensionality: $\text{PR}^{a}=1$. All dimensions are filled equally by activity.
An advantageous property of the participation ratio is that it can be expressed as the ratio of the squared-trace and Frobenius norm of $\bm{C}^{a}(0)$, both of which can be further expressed in terms of matrix elements $C_{ij}^{a}(0)$:
$\text{PR}^{a}$ $=\frac{\left(\text{tr}\bm{C}^{a}(0)\right)^{2}}{\lVert\bm{C}(0)\rVert_{F}^{2}}$
$=\frac{C^{a}(0)^{2}}{\frac{1}{N}\sum_{i=1}^{N}C_{ii}(0)^{2}+\frac{1}{N}\sum_{i\neq j}C_{ij}(0)^{2}},$ (21)
where $C^{a}(0)$ is given by Eq. 19. Assuming that diagonal elements are uniform with negligible fluctuations, i.e., $C^{a}_{ii}(0)=C^{a}(0)$ for all $i$ (as occurs, for example, under i.i.d. connectivity [10]), this simplifies to
$\text{PR}^{a}=\frac{C^{a}(0)^{2}}{C^{a}(0)^{2}+\psi^{a}(0,0)},$ (22)
where, following [30], we define the four-point function
$\psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i\neq j}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}),$ (23)
where $\bm{\tau}=(\tau_{1},\tau_{2})$. The three scaling behaviors of dimensionality outlined above correspond in the limit $N\to\infty$ to $\psi^{a}(0,0)\to\infty$}, $\psi^{a}(0,0)\to\text{const.}$}, or $\psi^{a}(0,0)\to 0$}, respectively. Since $\psi^{a}(0,0)$ is given by a sum over $\mathcal{O}(N^{2})$ squared cross-covariances times $1/N$, its magnitude is $N$ times that of an individual squared cross-covariance. Both i.i.d. matrices and the extensive-rank matrices generated by the random-mode model lead to $\mathcal{O}(1/\sqrt{N})$ cross-covariances. Thus, $\psi^{a}(0,0)\to\text{const.}$ as $N\to\infty$, leading to nontrivial extensive dimensionality, $0<\text{PR}^{a}<1$.
In this paper, we focus on a related function $\Psi^{a}(\bm{\tau})$ that includes the diagonal terms:
$\Psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i,j=1}^{N}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}).$ (24)
@@ III.3 Four-point functions, dimension of activity, and principal-component timescales
As per the duality between neuronal and temporal covariances, $\Psi^{a}(\bm{\tau})$ can also be expressed in terms of the time-by-time covariance $C^{a}(t_{1},t_{2})$ as
@@
The neuron-by-neuron (Eq. 24) and time-by-time (Eq. 25) definitions of $\Psi^{a}(\bm{\tau})$ form the basis of the cavity and path-integral calculations of this function, respectively. Finally, the dimension of activity is given by
$\text{PR}^{a}=\frac{C^{a}(0)^{2}}{\Psi^{a}(0,0)},$ (26)
which, unlike Eq. 22, holds even when single-neuron variances $C^{a}_{ii}(0)$ are nonuniform across neurons.
To study timescales of collective activity, we consider the principal components of activity:
$p^{a}_{k}(t)=\frac{1}{\sqrt{\lambda^{a}_{k}}}(\bm{v}^{a}_{k})^{T}\bm{a}(t).$ (27)
These principal components are the basis of much of modern analysis of high-dimensional neural data. They all have unit variance due to the $1/\sqrt{\lambda^{a}_{k}}$ normalization, but potentially very different characteristic timescales (Fig. 9a). To extract the timescales of just the leading components, we weight them by the squares of their corresponding eigenvalues. This gives: